Consider a solution initially containing 0.40 mol fluoride anion and 0.30 mol of hydrogen fluoride (HF). What is the pH after addition of 0.20 mol of HCl to this solution? (HF, Ka = 7.2 × 10(–4))? 



1
Expert's answer
2022-04-20T05:29:02-0400

F- + HCl = Cl- + HF


n(F-) = n(F-)initial - n(F-)reacted = 0.40 - 0.20 =0.20 mol

n(HF) = n(HF)initial + n(HF)produced in eaction = 0.30 + 0.20 = 0.50 mol


pH = pKa + log10"\dfrac{[F^-]}{[HF]}" = - log10(7.2*10-4) + log10(0.20/0.50) = 2.74